OpenStudy (anonymous):

find the height of an equilateral triangle whose area is equal to the area of a circle with circumference a 10cm

OpenStudy (anonymous):

Apple pI r squared ( so ) A=nr^2 (where n is pI) (A is Area) (r is radius) ( and ) Chicken pI dinner ( so ) C = nd (where n is pI) (C is Circumferance) (D is diameter (2xr)) use this to help

OpenStudy (anonymous):

10= 3.14... x D D/2 = R Area = 3.14... x R^2 Area = Equilateral - from there i dunno